Rowland & Sons Air Transport Service, Inc. has been in operation for three years. The following transactions occurred in February Feb. 1 Paid $555 for rent of hangar space in February Feb. 4 Received customer payment of $1,100 to ship several items to Philadelphia next month. Feb. 7 Flew cargo from Denver to Dallas; the customer paid in full ($1,100 cash). Feb. 10 Paid pilots $1,850 in wages for fying in February Feb. 14 Paid $164 for an advertisement run in the local paper on February 14 Feb. 18 Flew cargo for two customers from Dallas to Albuquerque for $2.365; one customer paid $825 cash and the other asked to be billed $1,540 Feb. 25 Purchased on account $1,570 in supplies for future use on the planes. Required: 1. Prepare accrual basis journal entries for each transaction. (If no entry is required for a transaction/event, select No Journal Entry Required in the first account field.) View transaction list Journal entry worksheet 6 Paid $553 for rent of hangar space in February. Record the transaction. Note: Enter debits before credits. Date General Journal Debit Credit Feb 01 Record entry Clear entry View general journal
2. Calculate the companys preliminary net income Preliminary Net Income 3. Calculate the companys net profit margin expressed as a percent Round your answer to 1 decimal place.) Net Profit Margin
Answer #1
1)
Feb-01 Rent expense 555
Cash 555
Feb-04 Cash 1000
Unearned freight revenue 1000
Feb-07 Cash 1100
Freight revenue 1100
Feb-10 Wages and salaries expense 1850
Cash 1850
Feb-14 Advertisement expenses 164
Cash 164
Feb-18 Cash 825
Freight revenue 825
Accounts receivable 2365
Freight revenue 2365
Feb-25 Supplies 1570
Accounts payable 1570
2) Net Income:
Freight revenue = 1100+825+2365 = 4290
Expenses = 555+1850+164 = 2569
Net income 1721
3) Net profit margin = 1721/4290 = 40.1%